4. Predict the structure of the compound that gives the spectrum shown below. Upon shaking the sample with D20, the doublet at 2.2 ppm disappears from the spectrum and the peak at 3.7 ppm simplifies to a clean sextet. Calculate the degree of unsaturation of the compound. What functional group must this compound contain? C4H10O 0.96 1 1.00 .98 3.08.08 7.5 7.0 6.6 5.0 4.5 2.5 2.0 1.5 0.0
Answer #1

degree of unsaturation = C - (H/2) +1 = 4 -5+1 = 0

Compound has no double bond

Most probable structure is CH3-CH(CH3) -CH2OH

peak around 2ppm is due to -CH2 attached to electronegative O atom. It is doublet.